A certain heat engine operating on a Carnot cycle
Answer #1

Q = 370 J

Th = 145 deg C ; eta = 24%

D)The netropy of the world changes by "zero"

delta S = 0 J/K

Since, Carnot cycle has reversible process.

E)we need to find W for this.

eta = W/Q

W = eta x Q = 0.24 x 370 = 88.8 J

PE = W

m g h = 88.8

m = 88.8/g h = 88.8/9.8 x 25 = 0.362 kg

Hence, m = 0.362 kg.